OpenStudy (anonymous):

tan58(degrees)=h/41........... i need to find the hight of the triangle

OpenStudy (raden):

cross multiply : h = 41 * tan58 = ... (use the calculator)

OpenStudy (anonymous):

yea im not sure how to do this stuff i need to get this right or i fail. can you find the answer for me

OpenStudy (anonymous):

\[\tan58^0= \frac{h}{41}\] \[1.6= \frac{h}{41} \rightarrow h= 41\times 1.6 =65.6\]

OpenStudy (anonymous):

@chrissaxe